OpenStudy (anonymous):

The price of a ticket for a certain concert is $5 for a child, $7 for a student, and $10 for an adult. Suppose that there were 260 people in attendance and that the total amount received in ticket sales was $1,932. Moreover, suppose that there were three times as many children at the concert as there were adults. How many children, students, and adults were at the concert? Find a system of linear equations which models this problem. Be sure to clearly identify the variables used. DO NOT SOLVE THE SYSTEM

OpenStudy (anonymous):

C + S + A = 260 C = 3A 5C + 7S + 10A = 1932

OpenStudy (anonymous):

is that the system of linear equations

OpenStudy (anonymous):

Yes : 3 vars, 3 equations. You can solve it.

OpenStudy (anonymous):

i dont know how haha, very lost

OpenStudy (anonymous):

Because C = 3A we have the following eq: S + 4 A = 260 25 A + 7 S =1932

OpenStudy (anonymous):

thats the complete answer

OpenStudy (anonymous):

No, that is the following step to resolve the system of 3 linear equations: reduce it to a system of 2 linear equations. Will you know the following step, to reduce it to a linear equation?

OpenStudy (anonymous):

to be honest i dont understand the question at all

OpenStudy (anonymous):

The good answer to your question is the following: C + S + A = 260 C = 3A 5C + 7S + 10A = 1932 with C for number of child, S for nb of students, and A for nb of adults.

OpenStudy (anonymous):

C number of child, S nb of students, and A nb of adults. C + S + A = 260 because there are 260 people. C = 3A because there are 3 times more child than adults. 5C + 7S + 10A = 1932 because the tickets sales was 1932$.